OpenStudy (anonymous):

2. What is a simpler form of the radical expression? square root of g to the sixth power (1 point)

OpenStudy (stamp):

do ur hw

OpenStudy (stamp):

\[\sqrt{g^6}=(g^6)^{1/2}\]

OpenStudy (stamp):

\[(x^a)^b=x^{ab}\]